OpenStudy (anonymous):

hi my name is pj and i need help with this problem? solve the quadratic equation x2+6x+8=0

OpenStudy (anonymous):

You can either factor it, or use the quadratic equation

OpenStudy (anonymous):

Which do you want to do?

OpenStudy (anonymous):

x = 1/6 (-8 - x2)

OpenStudy (anonymous):

{{x = -4}, {x = -2}} ignore my fist post

OpenStudy (anonymous):

how do i factor it?

OpenStudy (anonymous):

You look for two numbers that add together to give you 6 (the coefficient of the middle term), and multiply to give you 8 (the last term).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what two numbers add to make 6 and multiply to be 8 ?

OpenStudy (anonymous):

2 and 4

OpenStudy (anonymous):

Correct. So we can factor \(x^2+6x+8\) into \((x+4)(x+2)\). And if our original expression was equal to 0 then the factored form will also equal 0. \((x+4)(x+2) = 0 \implies x= -4 \text{ or } x = -2\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(x+6)(x-8)=0 is that x2+6x-8=0?